OpenStudy (anonymous):

k-10 over 2=negative7 help polpak please

OpenStudy (anonymous):

\[\frac{k-10}{2}= -7 \rightarrow k-10 = -7*2\]

OpenStudy (anonymous):

x2 on both sides, cancelling the '/2,' then -14 + 10

OpenStudy (anonymous):

Solve for k

OpenStudy (anonymous):

ya got my head hurting xplain step by step

OpenStudy (anonymous):

multiply both sides by 2 to get rid of the fraction you get \[k - 10 = -14\] Now add 10 to both sides to get \[k = -4\]

OpenStudy (anonymous):

multiply both sides by 2 to get rid of the fraction you get \[k - 10 = -14\] Now add 10 to both sides to get \[k = -4\]

OpenStudy (anonymous):

so k = -4

OpenStudy (anonymous):

yes that is correct